# CM points on Shimura curves via QM-equivariant isogeny volcanoes

### v2

@@ 7. Algebraic results on residue fields of CM points on $X^{D}(1)_{/\mathbb{Q}}$
Part (1) is [Cl23, Prop. 2.10], while part (2) is [CS23, Prop. 2.1] and part (3) follows from the combination of these two propositions. ∎
We now use Proposition 7.1 to get analogs of [Cl23, Prop. 2.10] and [CS23, Prop. 2.2], in which “rational ring class fields” are exchanged for those index $2$ subfields of rings class fields which arise as residue fields of CM points on $X^{D}(1)_{/\mathbb{Q}}$.
@@
follow by the same type of arguments as in the analogous case of rational ring class fields in [Cl23, Prop. 2.10] and [CS23, Prop. 2.2], using that $K(x)\cong K(x_{1})\cdot K(x_{2})\cong K(f_{1})\cdot K(f_{2})$ via Proposition 3.3 (note that the assumption that $x$ is fixed by $\sigma$ forces $f^{2}\Delta_{K}<-4$, so this proposition applies.).
@@ 8.1. Compiling across prime powers: $\Delta<-4$
[…]
@@ 8.2. Compiling across prime powers: $\Delta\in\{-3,-4\}$
Here, we determine how to compile residue field information across prime-power level for $\Delta\in\{-3,-4\}$. Our result here should be compared to [CS23, Prop. 82, Thm. 8.3], wherein more work is required due to the fact that residue fields of $-3$ and $-4$-CM points on $X_{0}(N)_{/\mathbb{Q}}$ do not always contain the CM field $K$.
@@ 8.3. The main algorithm
    • If $\Delta<-4$, then the count $c$ is given by the results of §8.1. If $\Delta\in\{-3,-4\}$, then this count is given by the results of §8.2 if $D>1$ and is given by the results of [CS23] if $D=1$.
@@ 8.4. Primitive residue fields of CM points on $X^{D}_{0}(N)_{/\mathbb{Q}}$
The proof follows exactly as in [Cl23, Thm. 9.2]; the main inputs here are the degrees of our residue fields, which are the same for our totally complex index $2$ subfields of ring class fields as they are for the rational ring class fields appearing in the $D=1$ modular curve study. ∎
@@ 9. CM points on $X_{1}^{D}(N)_{/\mathbb{Q}}$
In this section, we prove Theorem 1.2, showing that there is a very close relationship between CM points on the Shimura curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ and $X_{1}^{D}(N)_{/\mathbb{Q}}$. This is a generalization of [CS23, Thm. 1.2], which was specific to the $D=1$ case, and allows us to go from our understanding of the $\Delta$-CM locus on $X_{0}^{D}(N)_{/\mathbb{Q}}$ based on §8 to an understanding of, at the very least, degrees of CM points on $X_{1}^{D}(N)_{/\mathbb{Q}}$.
@@
We now assume that $x$ is a $\Delta$-CM point on $X_{0}^{D}(N)_{/\mathbb{Q}}$ with $\Delta\in\{-3,-4\}$ which is unramified with respect to the map to $X^{D}(1)_{/\mathbb{Q}}$. If $N=2$, then $\pi$ is an isomorphism, so the claim is trivial. If $N=3$, the fact mentioned above that there is one point lying over each elliptic point on $X^{D}(1)$ is exactly the inertness claim. For $N\geq 4$, we know that every point in $\pi^{-1}(x)$ is ramified with respect to the map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X^{D}(1)_{/\mathbb{Q}}$, giving the claimed ramification index. The residual degree is therefore at most the claimed residual degree in each case.
To provide the lower bound on the residual degree, we modify the argument of the $\Delta<-4$ case slightly in a predictable way. If $\Delta=-4$, then a representative for $x$ is well-defined up to quartic twist. We consider a representative of the form $(E_{1}\times E_{2},\iota,\mathcal{O}\cdot C)$ where $E_{1},E_{2}$ are $\mathfrak{o}(f)$-CM elliptic curves and $C\leq E_{1}[N]$ is a cyclic order $N$ subgroup (again, via the type of argument as in the $\Delta<-4$ case using Theorem 2.13). Let $q_{N}:\mathfrak{o}_{K}\rightarrow\mathfrak{o}_{K}/N\mathfrak{o}_{K}$ denote the quotient map. By tracking the action of Galois on a generator $P$ of $C$ we get a well-defined reduced mod $N$ Galois representation
@@ 10. Sporadic CM points on Shimura Curves
The natural map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ has non-trivial ramification exactly when $(D,N)$ is either Type I or Type II. In these cases, we have $d_{\text{CM}}(X_{0}^{D}(N))=2$, which is as small as possible as the $D>1$ assumption implies these curves have no rational points. The statements then follow immediately from the residual degrees with respect to this map provided by Theorem 1.2. ∎
@@ 10.3. Sporadic CM points
    There are at most $329$ pairs $(D,N)$, consisting of an indefinite quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ coprime to $D$, which do not appear among the $64$ listed in Table 1 and for which the Shimura curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ does not have a sporadic CM point. These are listed in the file unknown_sporadics.m in [Rep].
@@
    For all pairs $(D,N)$, consisting of an indefinite quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ coprime to $D$, which are not listed in Table 1 or among the $329$ referenced in the prior part (listed in unknown_sporadics.m) except possibly for $(91,5)$, the Shimura curve $X_{1}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point.
@@
For all of the $329$ pairs $(D,N)$ listed in unknown_sporadics.m, we have $d_{\text{CM}}(X_{0}^{D}(N))\in\{2,4,6\}$. For all but $56$ of these pairs, we have $d_{\textnormal{CM}}(X_{0}^{D}(N))=2$. For such pairs, it follows that the curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic (CM) point if and only if it is not bielliptic with a degree $2$ map to an elliptic curve over $\mathbb{Q}$ of positive rank. An extension of the results of [Rot02] mentioned in §10.2 to general level $N$ would then allow us to determine whether $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point for all but at most $56$ pairs $(D,N)$ with $D>1$. Such an extension will appear in work of the author and Oana Padurariu.

### v3

@@ 7. Algebraic results on residue fields of CM points on $X^{D}(1)_{/\mathbb{Q}}$
Part (1) is [Cl22, Prop. 2.10], while part (2) is [CS22, Prop. 2.1] and part (3) follows from the combination of these two propositions. ∎
We now use Proposition 7.1 to get analogs of [Cl22, Prop. 2.10] and [CS22, Prop. 2.2], in which “rational ring class fields” are exchanged for those index $2$ subfields of rings class fields which arise as residue fields of CM points on $X^{D}(1)_{/\mathbb{Q}}$.
@@
follow by the same type of arguments as in the analogous case of rational ring class fields in [Cl22, Prop. 2.10] and [CS22, Prop. 2.2], using that $K(x)\cong K(x_{1})\cdot K(x_{2})\cong K(f_{1})\cdot K(f_{2})$ via Proposition 3.3 (note that the assumption that $x$ is fixed by $\sigma$ forces $f^{2}\Delta_{K}<-4$, so this proposition applies.).
@@ 8.1. Compiling across prime powers: $\Delta<-4$
where $L\subsetneq K(\ell_{1}^{j_{1}}\cdots\ell_{r}^{j_{r}})$ is a totally complex, index $2$ subfield in the $s=0$ case. (Note that $\ell_{i}^{j_{i}}\Delta\in\{-12,-16,-27\}$ can only occur, due to the $\Delta<-4$ assumption, if $j_{i}=0$, so these possibilities do not require special attention here.)
@@ 8.2. Compiling across prime powers: $\Delta\in\{-3,-4\}$
Here, we determine how to compile residue field information across prime-power level for $\Delta\in\{-3,-4\}$. Our result here should be compared to [CS22, Prop. 82, Thm. 8.3], wherein more work is required due to the fact that residue fields of $-3$ and $-4$-CM points on $X_{0}(N)_{/\mathbb{Q}}$ do not always contain the CM field $K$.
@@ 8.3. The main algorithm
    • If $\Delta<-4$, then the count $c$ is given by the results of §8.1. If $\Delta\in\{-3,-4\}$, then this count is given by the results of §8.2 if $D>1$ and is given by the results of [CS22] if $D=1$.
@@ 8.4. Primitive residue fields of CM points on $X^{D}_{0}(N)_{/\mathbb{Q}}$
The proof follows exactly as in [Cl22, Thm. 9.2]; the main inputs here are the degrees of our residue fields, which are the same for our totally complex index $2$ subfields of ring class fields as they are for the rational ring class fields appearing in the $D=1$ modular curve study. ∎
@@ 9. CM points on $X_{1}^{D}(N)_{/\mathbb{Q}}$
In this section, we prove Theorem 1.2, showing that there is a very close relationship between CM points on the Shimura curves $X_{0}^{D}(N)_{/\mathbb{Q}}$ and $X_{1}^{D}(N)_{/\mathbb{Q}}$. This is a generalization of [CS22, Thm. 1.2], which was specific to the $D=1$ case, and allows us to go from our understanding of the $\Delta$-CM locus on $X_{0}^{D}(N)_{/\mathbb{Q}}$ based on §8 to an understanding of, at the very least, degrees of CM points on $X_{1}^{D}(N)_{/\mathbb{Q}}$.
@@
We now assume that $x$ is a $\Delta$-CM point on $X_{0}^{D}(N)_{/\mathbb{Q}}$ with $\Delta\in\{-3,-4\}$ which is unramified with respect to the map to $X^{D}(1)_{/\mathbb{Q}}$. If $N=2$, then $\pi$ is an isomorphism, so the claim is trivial. If $N=3$, the fact mentioned above that there is one point lying over each elliptic point on $X^{D}(1)$ is exactly the inertness claim. For $N\geq 4$, we know that every point in $\pi^{-1}(x)$ is ramified with respect to the map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X^{D}(1)_{/\mathbb{Q}}$, giving the claimed ramification index. The residue degree is therefore at most the claimed residue degree in each case.
To provide the lower bound on the residue degree, we modify the argument of the $\Delta<-4$ case slightly in a predictable way. If $\Delta=-4$, then a representative for $x$ is well-defined up to quartic twist. We consider a representative of the form $(E_{1}\times E_{2},\iota,\mathcal{O}\cdot C)$ where $E_{1},E_{2}$ are $\mathfrak{o}(f)$-CM elliptic curves and $C\leq E_{1}[N]$ is a cyclic order $N$ subgroup (again, via the type of argument as in the $\Delta<-4$ case using Theorem 2.13). Let $q_{N}:\mathfrak{o}_{K}\rightarrow\mathfrak{o}_{K}/N\mathfrak{o}_{K}$ denote the quotient map. By tracking the action of Galois on a generator $P$ of $C$ we get a well-defined reduced mod $N$ Galois representation
@@ 10. Sporadic CM points on Shimura Curves
The natural map $X_{1}^{D}(N)_{/\mathbb{Q}}\rightarrow X_{0}^{D}(N)_{/\mathbb{Q}}$ has non-trivial ramification exactly when $(D,N)$ is either Type I or Type II. In these cases, we have $d_{\text{CM}}(X_{0}^{D}(N))=2$, which is as small as possible as the $D>1$ assumption implies these curves have no rational points. The statements then follow immediately from the residue degrees with respect to this map provided by Theorem 1.2. ∎
@@ 10.3. Sporadic CM points
    There are at most $329$ pairs $(D,N)$, consisting of an indefinite quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ coprime to $D$, which do not appear among the $64$ listed in Table 1 and for which the Shimura curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ does not have a sporadic CM point. These are listed in Table LABEL:table:unknowns_table.
@@
    Let $(D,N)$ be a pair consisting of an indefinite quaternion discriminant $D>1$ over $\mathbb{Q}$ and a positive integer $N$ coprime to $D$. If $(D,N)$ is not listed in Table 1 or Table LABEL:table:unknowns_table and is not equal to $(91,5)$, then the Shimura curve $X_{1}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point.
@@
For all of the $329$ pairs $(D,N)$ listed in Table LABEL:table:unknowns_table, we have $d_{\text{CM}}(X_{0}^{D}(N))\in\{2,4,6\}$. For all but $56$ of these pairs, we have $d_{\textnormal{CM}}(X_{0}^{D}(N))=2$. For such pairs, it follows that the curve $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic (CM) point if and only if it is not bielliptic with a degree $2$ map to an elliptic curve over $\mathbb{Q}$ of positive rank. An extension of the results of [Rot02] mentioned in §10.2 to general level $N$ would then allow us to determine whether $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point for all but at most $56$ pairs $(D,N)$ with $D>1$. Such an extension will appear in work of the author and Oana Padurariu [PS24].
@@
Table 2. All $329$ pairs $(D,N)$ with $D>1$ for which we remain unsure whether $X_{0}^{D}(N)_{/\mathbb{Q}}$ has a sporadic CM point
$(6,17)$ $(6,23)$ $(6,25)$ $(6,35)$ $(6,41)$ $(6,43)$ $(6,47)$ $(6,49)$ $(6,53)$
$(6,55)$ $(6,59)$ $(6,61)$ $(6,65)$ $(6,67)$ $(6,71)$ $(6,73)$ $(6,77)$ $(6,79)$
$(6,83)$ $(6,85)$ $(6,89)$ $(6,91)$ $(6,95)$ $(6,97)$ $(6,101)$ $(6,103)$ $(6,107)$
$(6,109)$ $(6,113)$ $(6,115)$ $(6,119)$ $(6,121)$ $(6,125)$ $(6,127)$ $(6,131)$ $(6,133)$
$(6,137)$ $(6,139)$ $(6,143)$ $(6,145)$ $(6,149)$ $(6,151)$ $(6,155)$ $(6,157)$ $(6,161)$
$(6,163)$ $(6,167)$ $(6,169)$ $(6,173)$ $(6,179)$ $(6,181)$ $(6,191)$ $(6,193)$ $(6,197)$
$(6,199)$ $(6,203)$ $(6,287)$ $(6,295)$ $(6,319)$ $(10,9)$ $(10,13)$ $(10,17)$ $(10,19)$
$(10,21)$ $(10,27)$ $(10,29)$ $(10,31)$ $(10,33)$ $(10,37)$ $(10,39)$ $(10,41)$ $(10,43)$
$(10,47)$ $(10,49)$ $(10,51)$ $(10,53)$ $(10,57)$ $(10,59)$ $(10,61)$ $(10,63)$ $(10,67)$
$(10,69)$ $(10,71)$ $(10,73)$ $(10,77)$ $(10,79)$ $(10,83)$ $(10,87)$ $(10,89)$ $(10,91)$
$(10,97)$ $(10,103)$ $(10,119)$ $(10,141)$ $(10,161)$ $(10,191)$ $(14,3)$ $(14,9)$ $(14,11)$
$(14,13)$ $(14,15)$ $(14,17)$ $(14,19)$ $(14,23)$ $(14,25)$ $(14,27)$ $(14,29)$ $(14,31)$
$(14,33)$ $(14,37)$ $(14,39)$ $(14,41)$ $(14,43)$ $(14,47)$ $(14,53)$ $(14,59)$ $(14,61)$
$(14,87)$ $(14,95)$ $(15,4)$ $(15,7)$ $(15,8)$ $(15,11)$ $(15,13)$ $(15,14)$ $(15,16)$
$(15,17)$ $(15,19)$ $(15,22)$ $(15,23)$ $(15,26)$ $(15,28)$ $(15,29)$ $(15,31)$ $(15,32)$
$(15,34)$ $(15,37)$ $(15,41)$ $(15,43)$ $(15,47)$ $(21,2)$ $(21,4)$ $(21,5)$ $(21,8)$
$(21,10)$ $(21,11)$ $(21,13)$ $(21,16)$ $(21,17)$ $(21,19)$ $(21,23)$ $(21,25)$ $(21,29)$
$(21,31)$ $(21,38)$ $(22,7)$ $(22,9)$ $(22,13)$ $(22,15)$ $(22,17)$ $(22,19)$ $(22,21)$
$(22,23)$ $(22,25)$ $(22,27)$ $(22,29)$ $(22,31)$ $(22,35)$ $(22,37)$ $(22,51)$ $(26,3)$
$(26,5)$ $(26,7)$ $(26,9)$ $(26,11)$ $(26,15)$ $(26,17)$ $(26,19)$ $(26,21)$ $(26,23)$
$(26,25)$ $(26,29)$ $(26,31)$ $(33,2)$ $(33,4)$ $(33,5)$ $(33,7)$ $(33,8)$ $(33,10)$
$(33,13)$ $(33,16)$ $(33,17)$ $(33,19)$ $(34,3)$ $(34,5)$ $(34,7)$ $(34,9)$ $(34,11)$
$(34,13)$ $(34,15)$ $(34,19)$ $(34,23)$ $(34,29)$ $(34,35)$ $(35,2)$ $(35,3)$ $(35,4)$
$(35,6)$ $(35,8)$ $(35,9)$ $(35,11)$ $(35,12)$ $(35,13)$ $(38,3)$ $(38,5)$ $(38,7)$
$(38,9)$ $(38,11)$ $(38,13)$ $(38,17)$ $(38,21)$ $(39,4)$ $(39,5)$ $(39,7)$ $(39,8)$
$(39,10)$ $(39,11)$ $(39,31)$ $(46,3)$ $(46,5)$ $(46,7)$ $(46,9)$ $(46,11)$ $(46,13)$
$(46,15)$ $(46,17)$ $(51,2)$ $(51,4)$ $(51,5)$ $(51,7)$ $(51,8)$ $(51,10)$ $(51,11)$
$(51,20)$ $(55,2)$ $(55,3)$ $(55,4)$ $(55,7)$ $(55,8)$ $(57,2)$ $(57,4)$ $(57,5)$
$(57,7)$ $(58,3)$ $(58,5)$ $(58,7)$ $(58,9)$ $(58,11)$ $(58,13)$ $(62,3)$ $(62,5)$
$(62,7)$ $(62,9)$ $(62,11)$ $(62,15)$ $(65,2)$ $(65,3)$ $(65,4)$ $(65,7)$ $(69,2)$
$(69,4)$ $(69,5)$ $(69,7)$ $(69,11)$ $(74,3)$ $(74,5)$ $(74,7)$ $(77,2)$ $(77,3)$
$(77,4)$ $(77,5)$ $(77,6)$ $(82,3)$ $(82,5)$ $(82,7)$ $(86,3)$ $(86,5)$ $(86,7)$
$(87,2)$ $(87,4)$ $(87,5)$ $(87,8)$ $(94,3)$ $(94,5)$ $(94,7)$ $(95,2)$ $(95,3)$
$(106,3)$ $(106,5)$ $(106,7)$ $(111,2)$ $(111,4)$ $(118,3)$ $(118,5)$ $(119,2)$ $(119,3)$
$(119,6)$ $(122,3)$ $(122,5)$ $(122,7)$ $(129,2)$ $(129,7)$ $(134,3)$ $(134,5)$ $(134,9)$
$(143,2)$ $(143,4)$ $(146,3)$ $(146,7)$ $(159,2)$ $(166,3)$ $(183,5)$ $(194,3)$ $(215,2)$
$(215,3)$ $(326,3)$ $(327,2)$ $(335,2)$ $(390,7)$